# Infinity Synthetic Environments for Agentic Reinforcement Learning

### v2

@@ Abstract
Recent advances in large language model (LLM) have empowered autonomous agents to perform complex tasks that require multi-turn interactions with tools and environments. However, scaling such agent training is limited by the lack of diverse and reliable environments. In this paper, we propose Agent World Model (AWM), a fully synthetic environment generation pipeline. Using this pipeline, we scale to 1,000 environments covering everyday scenarios, in which agents can interact with rich toolsets (35 tools per environment on average) and obtain high-quality observations. Notably, these environments are code-driven and backed by databases, providing more reliable and consistent state transitions than environments simulated by LLMs. Moreover, they enable more efficient agent interaction compared with collecting trajectories from realistic environments. To demonstrate the effectiveness of this resource, we perform large-scale reinforcement learning for multi-turn tool-use agents. Thanks to the fully executable environments and accessible database states, we can also design reliable reward functions. Experiments on three benchmarks show that training exclusively in synthetic environments, rather than benchmark-specific ones, yields strong out-of-distribution generalization. The code is available at https://github.com/Snowflake-Labs/agent-world-model.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance in instruction following, reasoning, code generation, and tool-use (Chen et al., 2021; Schick et al., 2023; OpenAI, 2025; Anthropic, 2025a; Comanici et al., 2025; Guo et al., 2025). Agents powered by LLMs emerge as a promising paradigm for handling multi-step complex tasks in realistic environments (Nakano et al., 2021; Yao et al., 2023; Yang et al., 2024; Qin et al., 2024). However, training such agents often requires performing large-scale reinforcement learning (RL) on diverse environments that are relatively resource-scarce and expensive to scale.
Using real-world environments for training is prohibitively expensive and hard to scale, since many scenarios do not expose public APIs and RL training often requires agents to interact with them thousands of times in a stable and efficient manner (Dulac-Arnold et al., 2021; Qin et al., 2024; Xu et al., 2024a; Luo et al., 2025). Human-created environments are hard to scale and often lack diversity. For example, $\tau^{2}$-bench (Barres et al., 2025) and TheMCPCompany (Esfandiarpoor et al., 2025) only contain three and five environments, respectively, which is far from enough for training generic AI agents. Most existing synthetic research on agents focuses on task synthesis (Wang et al., 2023; Chen et al., 2024; Xie et al., 2025; Patil et al., 2024; Wang et al., 2026) and trajectory collection (Xu et al., 2024b; Li et al., 2025a; Song et al., 2024) rather than environment synthesis. Another line of research simulates the tool response or even the environment, where each state transition is generated by LLMs (Liu et al., 2024b; Lu et al., 2025; Li et al., 2025b, c; Chen et al., 2025). However, this approach is not reliable or efficient due to the hallucination issue (Wang et al., 2024; Kalai et al., 2025) and LLM’s high inference cost.
@@
To address this, we propose Agent World Model (AWM), an open-source pipeline that synthesizes executable tool-use environments at scale. The key insight is that agent environments share a common structure: a stateful backend, a tools interface layer, and task-specific success criteria. By decomposing synthesis into these three components, we can leverage LLMs to generate each part systematically while maintaining consistency. Our AWM mirrors how software is built in practice. Starting from a high-level scenario description (e.g., “an online shopping platform”), we first generate common user requirements (i.e., tasks) that users are likely to perform in this scenario. Then, we generate the database schema to define what entities and relations exist to fulfill these user requirements. This schema can guide the design of exposed interfaces (toolset) and help generate the backend code for the interfaces, ensuring each tool has a clear data model to operate on. The interface is exposed via Model Context Protocol (MCP) (Anthropic, 2024) for unified agent tool interaction with the environment. Finally, we generate verification code that compares the database state before and after agent execution, which augments an LLM-as-a-Judge to provide robust reward signals for RL training. Critically, each stage includes automated execution and simple self-correction: if generated code fails to run, we feed the error information back to the LLM for correction.
This pipeline enables us to scale to 1,000 unique environments spanning most real-world scenarios such as shopping, social media, finance, and travel. Each environment provides an executable sandbox where agents can interact with dozens of tools, and they fully support parallel isolated instances and are easy to reset or restart, which are important for efficient online RL. To validate AWM, we perform large-scale RL (each step with 1,024 environment instances) to train agents to use MCP tools to complete the task. In contrast to some works performing training on test environments, our environments and tasks are not tailored for any benchmarks or specific scenarios. Experimental results on three tool-use benchmarks demonstrate the generalization performance of our framework. In summary, our contributions are three-fold: (1) AWM, an open-source pipeline for automatic generation of executable tool-use environments with database-backed state consistency, (2) 1,000 ready-to-use diverse environments suitable for large-scale RL training, and (3) empirical results show that agents trained on AWM generalize to out-of-distribution environments.
@@ 2 Related Work
Tool-use Agents. Recent work shows that LLMs can use external tools to solve complex tasks (Qin et al., 2024; OpenAI, 2025; DeepSeek-AI et al., 2025; Team et al., 2025). Toolformer (Schick et al., 2023) trains tool-use LLMs by supervised learning. ToolLLM (Qin et al., 2024) curates real-world APIs and trains on LLM-generated trajectories, but uses simulated responses instead of tool execution. Gorilla (Patil et al., 2024) fine-tunes with API documentation to improve tool-use accuracy. ReAct (Yao et al., 2023) and SWE-agent (Yang et al., 2024) alternate reasoning and acting in interactive environments. However, most training data remains static or comes from small-scale environments. Existing benchmarks (Yao et al., 2024; Barres et al., 2025; Esfandiarpoor et al., 2025; Luo et al., 2025; Fan et al., 2025; Mo et al., 2025) either use real-world APIs or provide small-scale environments. This makes them hard to use as large-scale RL training grounds. The gap is the lack of diverse and executable environments that are efficient enough for RL. Such training needs extensive agent interactions, and it benefits from fast interactions and reliable state transitions.
@@
Environment Synthesis. As agentic RL grows, the need for diverse and executable environments becomes more urgent. Environment synthesis largely follows two directions: LLM-based simulation (Wang et al., 2024; Li et al., 2025c; Chen et al., 2025; Li et al., 2025b) and programming-based synthesis (Tang et al., 2024; Sullivan et al., 2025; Cai et al., 2025; Zhang et al., 2025a; Song et al., 2026). LLM-based simulation often simulates the environment by prompting reasoning model to generate state transition and observation. However, it suffers from hallucinations in state transition (Kalai et al., 2025; Wang et al., 2024). It is also expensive and inefficient for RL, since each environment step may require an LLM call. Programming-based synthesis builds environments through programming and database, where each state transition and observation are driven by the code. Some methods extract tool graphs from documentation and sample tool call sequences to build environments (Fang et al., 2025; Cai et al., 2025). Procedural generation has also been explored, for example with human designed type systems (Sullivan et al., 2025). AutoEnv (Zhang et al., 2025a) creates 36 game-like environments (e.g., maze navigation) for simulating heterogeneous worlds. EnvScaler (Song et al., 2026) is a concurrent work that synthesizes 191 interactive environments via code generation given existing task sets.
The proposed AWM differs in three aspects: (1) We synthesize environments from scratch without assuming predefined task sets or API documentation. This avoids human priors and enables scalable synthesis. (2) We use database-backed state management to enforce consistency and enable code-augmented verification for RL. (3) With 1,000 environments, 35,062 tools and 10,000 tasks paired with verification code synthesized, to the best of our knowledge, AWM is the largest open-source tool-use environment set to date.
@@ 3.1 Scenario Generation
We leverage the vast world knowledge of modern LLMs to generate diverse scenario descriptions (websites, apps, or common tools collection), seeded with 100 popular domain names. Unlike general web agents that navigate static content sites (e.g., news, wikis), we focus on stateful applications (e.g., e-commerce, CRM, management) that necessitate database interactions instead of information retrieval. To ensure quality, we employ a filtering pipeline: an LLM-based classifier selects scenarios involving core CRUD operations (create, read, update, delete), while embedding-based deduplication ensures diversity. We also cap over-represented categories to prevent the collection from collapsing to a few dominant types. This process yields 1,000 unique scenarios spanning finance, travel, retail, social media, and more, as shown in Figure 7 and Table 9.
@@ 3.3.1 Environment
Execution-based Self-Correction. Across all synthesis steps described above, we employ a simple self-correction mechanism to handle generation errors. After generating each component of the environment, we attempt to run it in an isolated environment and test the functionality. If any error occurs (e.g., runtime exceptions), we capture the error message and feed it back to the LLM together with the problematic code snippet, prompting the LLM to regenerate a corrected version. This process is repeated for up to five iterations or until the component executes successfully. In practice, we often find that this lightweight retry strategy is effective in repairing generated code, without requiring a more complex correction mechanism.
@@ 3.3.2 Pipeline Results and Analysis
Table 2: Environment complexity statistics. Agent steps and unique tools are measured using Claude-4.5-Sonnet (Anthropic, 2025a) as the agent backbone, with a maximum step budget of 20. About $13.7\%$ of tasks exceeded this budget.
@@
Table 3: Comparison with existing programming-based environments. Our AWM generates environments with minimal reliance (only a seed set of scenario names) on human or real APIs, achieving the largest scale with SQL-backed state consistency.
@@ 4.2 History-Aware Training
When deploying agents, history context is often managed by a dedicated framework that strategically truncates long interaction histories to avoid attention sink and improve efficiency (Liu et al., 2024a; Anthropic, 2025b; Zhang et al., 2025b). However, existing RL training pipelines may optimize policies using full histories, creating a distribution mismatch issue between training and inference.
@@ 5.1 Experimental Setup
Benchmarks. We evaluate agents on three tool-use and MCP benchmarks to assess out-of-distribution generalization: (1) The verified version of $\tau^{2}$-bench (Barres et al., 2025; Cuadron et al., 2025), which consists of multi-turn conversational agentic tasks across three representative scenarios: airline, retail, and telecom; (2) BFCLv3 (Patil et al., 2025), a comprehensive benchmark for function-calling ability evaluation, covering single-turn, multi-turn (long-context), synthetic tools, real-world tools, and hallucination tests, resulting in four evaluation categories: non-live, live, multi-turn, and hallucination; (3) MCP-Universe (Luo et al., 2025), a collection of real-world MCP servers spanning location navigation, financial analysis, browser automation, web search, and multi-server workflows. We exclude 3D design tasks that require GUI and repository management tasks that require authenticated access to GitHub or Notion.
Baselines. We compare against the following baselines: (1) Base: the original LLMs equipped with reasoning and tool-use capabilities, without additional training; (2) Simulator (Li et al., 2025b; Chen et al., 2025): agents trained with RL in LLM-simulated environments, where GPT-5 (OpenAI, 2025) serves as the environment transition model, using the same tasks and toolsets as AWM, to highlight the advantage of executable environments over simulated ones; (3) EnvScaler (Song et al., 2026): a concurrent method that synthesizes 191 programming-based environments for SFT and RL, included to compare the synthesis quality.
Implementation Details. The AWM pipeline is instantiated using GPT-5 (OpenAI, 2025), which is also used as the code-augmented LLM-as-a-Judge for task verification. We select Qwen3 thinking models (Yang et al., 2025) across 4B, 8B, and 14B scales as base agents due to their popular community adoption. Multi-turn RL training is implemented on top of AgentFly (Wang et al., 2025a) and verl (Sheng et al., 2024). Due to limited computation budget, we train agents on a subset of AWM consisting of 526 environments and 3,315 tasks. Each model is trained for up to 96 optimization steps with a fixed learning rate of $7\times 10^{-7}$. We use a batch size of 64 and 16 rollouts, resulting in 1,024 isolated environment instances launched per step. The sliding window size and maximum interaction turn are set to $w=3$ and $20$, respectively. More details are in Appendix A.
@@ 6.1 Quality of Synthesized Environments
We evaluate synthesized environments based on two critical criteria for scalable agentic RL: quality (tasks are executable with coherent schema and tool interfaces) and diversity (training does not collapse to near-duplicate environments).
Quality. Table 5 reports LLM-as-a-Judge scores on Task Feasibility (whether tasks are executable within the environment), Data Alignment (whether the data schema is coherent with the task), and Toolset Completeness (whether the toolset is complete and usable), where AWM consistently outperforms EnvScaler, indicating stronger end-to-end consistency from tasks $\rightarrow$ database $\rightarrow$ interface. Despite our environments containing roughly 3x more code than those generated by EnvScaler (see Table 3), this increase in scale results in only a moderate rise in bugs, highlighting effective scalability. Both methods exhibit implementation issues at this scale, manual inspection on our environments attributes 44% of bugs to not handling edge input cases and 14% to operations conflicting with database constraints. During RL training, the environment error rate consistently remains low, around 4%, indicating overall reliability and usability of our environments. AWM yields fewer blocked tasks than EnvScaler, which is critical for RL because blocked tasks truncate exploration and inject systematically incorrect negative signals. To mitigate such imperfect environment issue, we design Environment Error reward in Sec. 3.3.1.
Diversity. Figure 3 analyzes diversity from semantic and topical perspectives. Embedding diversity remains stable as pool size grows, suggesting newly generated environments continue to add novel content rather than forming duplicates. Meanwhile, category coverage steadily increases, showing that AWM globally expands into new regions instead of collapsing to a few dominant domains. Overall, quality and diversity results suggest that AWM can scale to thousands of high-quality environments while maintaining diversity.

### v3

@@ Abstract
Recent advances in large language model (LLM) have empowered autonomous agents to perform multi-turn interactions with tools and environments. However, scaling such agent training is limited by the lack of diverse and reliable environments. In this paper, we propose Agent World Model (AWM), a fully synthetic environment generation pipeline. Using this pipeline, we scale to 1,000 environments covering everyday scenarios, in which agents can interact with rich toolsets and obtain high-quality observations. Notably, these environments are code-driven and backed by databases, providing more reliable and consistent state transitions than environments simulated by LLMs. Moreover, they enable more efficient agent interaction compared with collecting trajectories from realistic environments. To demonstrate the effectiveness of this resource, we perform large-scale reinforcement learning for multi-turn tool-use agents. Thanks to the fully executable environments and accessible database states, we can also design reliable reward functions. Experiments on three benchmarks show that training exclusively in synthetic environments, rather than benchmark-specific ones, yields strong out-of-distribution generalization. The code is available at https://github.com/Snowflake-Labs/agent-world-model.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance in instruction following, reasoning, code generation, and tool-use (Chen et al., 2021; Schick et al., 2023; OpenAI, 2025; Anthropic, 2025b; Comanici et al., 2025; Guo et al., 2025). Agents powered by LLMs emerge as a promising paradigm for handling multi-step complex tasks in realistic environments (Nakano et al., 2021; Yao et al., 2023; Yang et al., 2024; Qin et al., 2024). However, training such agents often requires performing large-scale reinforcement learning (RL) on diverse environments that are relatively resource-scarce and expensive to scale.
@@
Using real-world environments for training is prohibitively expensive and hard to scale, since many scenarios do not expose public APIs and RL training often requires agents to interact with them thousands of times in a stable and efficient manner (Dulac-Arnold et al., 2021; Qin et al., 2024; Xu et al., 2024a; Luo et al., 2025b). Human-created environments are hard to scale and often lack diversity. For example, $\tau^{2}$-bench (Barres et al., 2025) and TheMCPCompany (Esfandiarpoor et al., 2025) only contain three and five environments, respectively, which is far from enough for training generic AI agents. Most existing synthetic research on agents focuses on task synthesis (Wang et al., 2023; Chen et al., 2024; Xie et al., 2025; Patil et al., 2024; Wang et al., 2026) and trajectory collection (Xu et al., 2024b; Li et al., 2025a; Song et al., 2024) rather than environment synthesis. Another line of research simulates the tool response or even the environment, where each state transition is generated by LLMs (Liu et al., 2024b; Lu et al., 2025; Li et al., 2025c, b; Chen et al., 2025). However, this approach is not reliable or efficient due to the hallucination issue (Wang et al., 2024; Kalai et al., 2025) and LLM’s high inference cost.
@@
To address this, we propose Agent World Model (AWM), an open-source pipeline that synthesizes executable tool-use environments at scale. AWM is analogous to learned world models in model-based RL but realized via code-driven environments rather than neural dynamics. By decomposing synthesis into these three components, we can leverage LLMs to generate each part systematically while maintaining consistency. Our AWM mirrors how software is built in practice. Starting from a high-level scenario description (e.g., “an online shopping platform”), we first generate common user requirements (i.e., tasks) that users are likely to perform in this scenario. Then, we generate the database schema to define what entities and relations exist to fulfill these user requirements. This schema can guide the design of exposed interfaces (toolset) and help generate the backend code for the interfaces, ensuring each tool has a clear data model to operate on. The interface is exposed via Model Context Protocol (MCP) (Anthropic, 2024) for unified agent tool interaction with the environment. Finally, we generate verification code that compares the database state before and after agent execution, which augments an LLM-as-a-Judge to provide robust reward signals for RL training. Critically, each stage includes automated execution and simple self-correction: if generated code fails to run, we feed the error information back to the LLM for correction.
This pipeline enables us to scale to 1,000 unique environments spanning most real-world scenarios such as shopping, social media, finance, and travel. Each environment provides an executable sandbox where agents can interact with dozens of tools, and they fully support parallel isolated instances and are easy to reset or restart, which are important for efficient online RL. To validate AWM, we perform large-scale RL (each step with 1,024 environment instances) to train agents to use MCP tools to complete the task. In contrast to recent work that trains agents using benchmark-specific environments or simulators conditioned on benchmark task sets (Li et al., 2025c; Chen et al., 2025; Luo et al., 2025a; Prabhakar et al., 2025; Wang et al., 2025b), our environments and tasks are not tailored for any benchmarks or specific scenarios. Experimental results on three tool-use benchmarks demonstrate the generalization performance of our framework. In summary, our contributions are three-fold: (1) AWM, an open-source pipeline for automatic generation of executable tool-use environments with database-backed state consistency, (2) 1,000 ready-to-use diverse environments suitable for large-scale RL training, and (3) empirical results show that agents trained on AWM generalize to out-of-distribution unseen environments.
@@ 2 Related Work
Tool-use Agents. Recent work shows that LLMs can use external tools to solve complex tasks (Qin et al., 2024; OpenAI, 2025; DeepSeek-AI et al., 2025; Team et al., 2025). Toolformer (Schick et al., 2023) trains tool-use LLMs by supervised learning. ToolLLM (Qin et al., 2024) curates real-world APIs and trains on LLM-generated trajectories, but uses simulated responses instead of tool execution. Gorilla (Patil et al., 2024) fine-tunes with API documentation to improve tool-use accuracy. ReAct (Yao et al., 2023) and SWE-agent (Yang et al., 2024) alternate reasoning and acting in interactive environments. However, most training data remains static or comes from small-scale environments. Existing benchmarks (Yao et al., 2024; Barres et al., 2025; Esfandiarpoor et al., 2025; Luo et al., 2025b; Fan et al., 2025; Mo et al., 2025) either use real-world APIs or provide small-scale environments. This makes them hard to use as large-scale RL training grounds. The gap is the lack of diverse and executable environments that are efficient enough for RL which needs extensive agent interactions, and it benefits from fast interactions and reliable state transitions.
@@
Environment Synthesis. As agentic RL grows, the need for diverse and executable environments becomes more urgent. Environment synthesis largely follows two directions: LLM-based simulation (Wang et al., 2024; Li et al., 2025b; Chen et al., 2025; Li et al., 2025c) and programming-based synthesis (Tang et al., 2024; Sullivan et al., 2025; Cai et al., 2025; Zhang et al., 2025a; Song et al., 2026). LLM-based simulation often simulates the environment by prompting reasoning model to generate state transition and observation. However, it suffers from hallucinations in state transition (Kalai et al., 2025; Wang et al., 2024). It is also expensive and inefficient for RL, since each environment step may require an LLM call. Programming-based synthesis builds environments through programming and database, where each state transition and observation are driven by the code. Some methods extract tool graphs from documentation and sample tool call sequences to build environments (Fang et al., 2025; Cai et al., 2025). Procedural generation has also been explored, for example with human designed type systems (Sullivan et al., 2025). AutoEnv (Zhang et al., 2025a) creates 36 game-like environments (e.g., maze navigation) for simulating heterogeneous worlds. EnvScaler (Song et al., 2026) is a concurrent work that synthesizes 191 interactive environments via code generation given existing task sets.
AWM differs in three aspects: (1) We synthesize from scratch without assuming predefined task sets or API documentation, which avoids human priors. (2) We use database-backed state management to enforce consistency and enable code-augmented verification for RL. (3) With 1,000 environments, 35,062 tools and 10,000 tasks paired with verification code synthesized, to the best of our knowledge, AWM is the largest open-source environment set to date.
@@ 3.1 Scenario Generation
We leverage the vast world knowledge of modern LLMs to generate diverse scenario descriptions (websites, apps, or common tools collection), seeded with 100 popular domain names. Unlike general web agents that navigate static content sites (e.g., news, wikis), we focus on stateful applications (e.g., e-commerce, CRM, management) that necessitate database interactions instead of information retrieval. To ensure quality, we employ a filtering pipeline: an LLM-based classifier selects scenarios involving core CRUD operations (create, read, update, delete), while embedding-based deduplication ensures diversity. We also cap over-represented categories to prevent the collection from collapsing to a few dominant types. This process yields 1,000 unique scenarios spanning finance, travel, retail, social media, workflow, user management, and more, as shown in Figure 7 and Table 15 in Appendix A.1.
@@ 3.3.1 Environment
Execution-based Self-Correction. Across all synthesis steps described above, we employ a simple self-correction mechanism to handle generation errors. After generating each component of the environment, we attempt to run it in an isolated environment and test the functionality. If any error occurs (e.g., runtime exceptions), we capture the error message and feed it back to the LLM together with the problematic code snippet, prompting the LLM to regenerate a corrected version. This process is repeated for up to five iterations or until the component executes successfully. In practice, we often find that this lightweight retry strategy is effective in repairing generated code, without requiring a more complex correction mechanism. During the synthesis process, we finally allow up to $10\%$ errors in each stage to reduce costs, while acknowledging that setting a stricter threshold or using modern coding agent harness such as Claude Code can further improve the quality of the environments at the cost of more correction iterations.
@@ 3.3.2 Pipeline Results and Analysis
Table 2: Environment complexity statistics. Agent steps and unique tools are measured using Claude-4.5-Sonnet (Anthropic, 2025b) as the agent backbone, with a step budget of 20. The completion rate is $62.6\%$, and $13.7\%$ of tasks exceeded the step budget.
@@
Table 3: Comparison with existing programming-based environments. AWM generates environments with minimal reliance (only a seed set of scenario names) on human or real APIs, achieving the largest scale with SQL-backed state consistency. A direct empirical comparison with AutoForge is not feasible because its code and environments are not released.
@@ 4.2 History-Aware Training
When deploying agents, history context is often managed by a dedicated framework that strategically truncates long interaction histories to avoid attention sink and improve efficiency (Liu et al., 2024a; Anthropic, 2025a; Zhang et al., 2025b). However, existing RL training pipelines may optimize policies using full histories, creating a distribution mismatch issue between training and inference.
@@ 5.1 Experimental Setup
Benchmarks. We evaluate agents on three tool-use and MCP benchmarks to assess out-of-distribution generalization: (1) The verified version of $\tau^{2}$-bench (Barres et al., 2025; Cuadron et al., 2025), which consists of multi-turn conversational agentic tasks across three representative scenarios: airline, retail, and telecom; (2) BFCLv3 (Patil et al., 2025), a comprehensive benchmark for function-calling ability evaluation, covering single-turn, multi-turn (long-context), synthetic tools, real-world tools, and hallucination tests, resulting in four evaluation categories: non-live, live, multi-turn, and hallucination; (3) MCP-Universe (Luo et al., 2025b), a collection of real-world MCP servers spanning location navigation, financial analysis, browser automation, web search, and multi-server workflows. We exclude 3D design tasks that require GUI and repository management tasks that require authenticated access to GitHub or Notion.
Baselines. We compare against the following baselines: (1) Base: the original LLMs equipped with reasoning and tool-use capabilities, without additional training; (2) Simulator (Li et al., 2025c; Chen et al., 2025): agents trained with RL in LLM-simulated environments, where GPT-5 (OpenAI, 2025) serves as the environment transition model, using the same tasks and toolsets as AWM, to highlight the advantage of executable environments over simulated ones; (3) EnvScaler (Song et al., 2026): a concurrent method that synthesizes 191 programming-based environments for SFT and RL, included to compare the synthesis quality.
Implementation Details. The AWM pipeline is instantiated using GPT-5 (OpenAI, 2025), which is also used as the code-augmented LLM-as-a-Judge for task verification. We select Qwen3 thinking models (Yang et al., 2025) across 4B, 8B, and 14B scales as base agents due to their popular community adoption. Multi-turn RL training is implemented on top of AgentFly (Wang et al., 2025a) and verl (Sheng et al., 2024). Due to limited computation budget, we train agents on a subset of AWM consisting of 526 environments and 3,315 tasks. Each model is trained for up to 96 optimization steps with a fixed learning rate of $7\times 10^{-7}$. We use a batch size of 64 and 16 rollouts, resulting in 1,024 isolated environment instances launched per step. The sliding window size and maximum interaction turn are set to $w=3$ and $20$, respectively. For evaluation, we adapt benchmarks to MCP interface by requiring agents to use tools via two unified tools: list_tools and call_tool. More implementation details can be found in Appendix A.
@@ 5.2 Main Results
Gains across complexity. We stratify BFCLv3 and $\tau^{2}$-bench tasks by inherent difficulty (number of required tool calls / ground-truth actions) in Table 5. AWM improves over Base across all complexity buckets on both benchmarks. The absolute gain shrinks on harder tasks while the relative gain remains substantial, which indicates the agent’s performance on hard tasks relies on its intrinsic capabilities.
Table 5: Benchmark task complexity stratification with 8B agent. MCP-Universe does not have ground truth trajectory reference or other complexity indicators, thus being excluded.
BFCLv3 (by # tool calls) $\tau^{2}$ (by # GT actions)
Bucket Base AWM $\Delta$ Base AWM $\Delta$
Simple 53.6 80.3 +26.7 32.7 41.9 +9.3
Medium 60.0 75.3 +15.3 22.7 28.8 +6.2
Hard 43.9 45.0 +1.1 20.5 25.0 +4.5
@@ 6.1 Quality of Synthesized Environments
Table 7: Quality analysis of AWM synthesis pipeline on 100 environments (1,000 tasks) under three different generator models.
Metric Generator GPT-5 Claude-4.5 Qwen3.5
Pipeline Success Rate (%, $\uparrow$)
Database Synthesis 88.3 100.0 79.0
Sample Data Synthesis 88.2 100.0 97.0
Env. Code Synthesis 86.8 99.0 77.0
Quality Score (1–5, $\uparrow$)
Task Feasibility 3.99 3.84 3.15
Data Alignment 4.84 4.92 4.32
Toolset Completeness 4.98 4.97 4.30
Bug Statistics ($\downarrow$)
Envs w/ Bugs (%) 83 89 100
# Bugs per Env 2.70 2.11 3.82
Diversity ($\uparrow$)
Mean Pairwise Distance 0.34 0.31 0.35
Table 8: Complexity-stratified Pass@1 (%) evaluation on synthesized tasks by GPT-5.1 and Claude-4.5-Sonnet. Buckets are defined by required tool calls: 1-3, 4-6, 7-10 and 11+.
Metric Bucket Simple Medium Hard Very Hard
Tasks Ratio (%) 35.1 31.5 15.7 17.7
Avg. Steps 3.3 5.8 9.1 17.9
GPT-5.1 61.7 27.0 11.9 3.0
Sonnet-4.5 68.4 71.8 63.3 31.0
Both Pass 45.3 24.9 11.5 3.0
Neither Pass 15.2 26.2 36.3 69.0
We evaluate synthesized environments along three axes that are crucial for agentic RL training: quality (tasks are executable with coherent environment), difficulty (synthesized tasks bring new information), and diversity (training does not collapse to near-duplicate environments).
@@
Quality. Table 6 reports LLM-as-a-Judge scores on Task Feasibility, Data Alignment, and Toolset Completeness, where AWM consistently outperforms EnvScaler under both GPT-5.1 and Claude-4.5-Sonnet judges, indicating stronger end-to-end consistency from tasks $\rightarrow$ database $\rightarrow$ interface. Despite our environments containing roughly 3x more code than those generated by EnvScaler (Table 3), this increase in scale results in only a moderate rise in bugs, highlighting effective scalability. Both methods exhibit implementation issues at this scale: manual inspection on our environments attributes 44% of bugs to unhandled edge input cases (e.g., missing null/boundary validation) and 14% to operations conflicting with database constraints (e.g., foreign-key or uniqueness violations), which together account for the majority of bugs and are typical of auto-generated code at this scale. During RL training, the environment error rate consistently remains low, around 4%, most of which are internal errors caused by specific agent behaviors, indicating overall reliability and usability of our environments. AWM yields fewer blocked tasks than EnvScaler, which is critical for RL because blocked tasks truncate exploration and inject systematically incorrect negative signals. To mitigate such imperfect-environment issues, we design the Environment Error reward in Sec. 3.3.1.
To verify the universality of our pipeline, we further test Claude-4.5-Sonnet and open-source Qwen3.5-122B-A10B for the synthesis in Table 7. The pipeline is largely model-agnostic: Claude-4.5 attains 99% environment-code success and matches GPT-5 on quality, while Qwen3.5 still reaches 77% with usable but bug-heavier environments. Diversity is essentially constant across generators, confirming that it comes from the pipeline design rather than the used model.
Difficulty. We present a complexity-stratified analysis of AWM tasks in Table 8. Overall, the two advanced models achieve only 36.1% and 62.6% Pass@1 respectively, confirming that AWM tasks are non-trivial (GPT-5.1’s lower rate may be partly due to its occasional refusal to follow the system prompt). Pass rates decrease monotonically with complexity for both models, and 69% of Very-Hard tasks are unsolved by either. Furthermore, from our RL training logs, 27.1%, 23.7%, and 28.2% of tasks are never solved even once by Qwen3-4B/8B/14B throughout the entire training process (14B only optimizes for 32 steps), providing further evidence that AWM generates genuinely challenging tasks that remain difficult even after extensive RL exploration.
Diversity. Figure 3 analyzes diversity from semantic and topical perspectives. Embedding diversity remains stable as pool size grows, suggesting newly generated environments continue to add novel content rather than forming duplicates. Meanwhile, category coverage steadily increases, showing that AWM globally expands into new regions instead of collapsing to a few dominant domains. We further conduct a code-level similarity study using AST-based duplicate detection and token/n-gram Jaccard metrics. The cross-environment AST function duplicate rate is 0.0% and endpoint name Jaccard is 0.004 (Appendix B.3), confirming that diversity holds even at the implementation level.